Previous

Calendar

October

2007

Next
Sunday Monday Tuesday Wednesday Thursday Friday Saturday

30

JD273
First significant snowfall (on Sat night).

1

JD274

2

JD275
More snow (Mon night) + strong winds.

3

JD276
Site Visit (JB, MP, AW, MW, JM, LL): Details

removed photo diode wands

4

JD277
Site Visit (JB, MP, SpB, NT, SL): Details , Photos (17)

reboot of russter2 ~10:00 MDT

swapped four flasks, done=2; switched robot program to semi-winter mode (seasonal flag=1, collects 4 flasks every week)

moved ATI 2-d sonic from 20cm to tower @ ~9m

from 13:40-14:10 MST: Finished blow-testing the HYDRA inlets

5

JD278

6

JD279

7

JD280

8

JD281
Site Visit (SO+): Details

Columbus Day

9

JD282
Site Visit (JB, SpB, MP, NT): Details , Photos (11)

from 9:30-10:00 MST: hydra gast pump off, switched around inlet connections (see details)

at ~9:30 MST, added KZ PAR-LITE sensor as PAR2.up to CNR logger

at 11:30 MST, removed a pair of bugs stuck in the rain gauge

from 11:17-11:55 MST, replaced 2 micron Nupro filters for the LI-6262 and re-zeroed the LI-6262

from 13:30-13:50 MST, fixed up bottom three thermocouples

downloaded from cnr, prof, met, nctc, soil, tcs, ucan, and ucb

10

JD283
Site Visit (JB, SpB, MW, SO+): Details , Photos (5)

at ~12:30 MST: HYDRA data system shuts down.

at 13:30 MST, NCAR TC cr23x logger stops working

at around 12:00 MST, reconnected refurbished 16m prop-vane

started hauling TRAM gear to new transect

11

JD284
Site Visit (JB, SpB, NT, AP, AT, SO+): Details

at 22:30 MST: HYDRA data system back up.

swapped four flasks

hauled more TRAM gear

12

JD285
Site Visit (JB, SpB, RaM, SO+): Details , Photos (6)

at 12:45 MST: HYDRA shutdown and removed from site.

removed NCAR TC cr23x logger to send back for repairs.

swapped mag. perch. trap and indicator tube for the 6251 @ ~12:20 MDT

all TRAM towers erected and guyed at the new transect

13

JD286

14

JD287
Heavy snowfall.

15

JD288
MP and AW take the gray Jeep to the PJ woodlands in southern CO to collect more caterpillars

16

JD289
Site Visit (JB, SpB, LL, NT, C, DB, JM): Details , Photos (19)

removed the KZ PAR-LITE sensor (PAR2.up) from the CNR logger

@ ~15:00 MDT we changed the nupro filter on the LI-6251 and swapped the mag. perch. trap and indicator tube

17

JD290
Ameriflux Meeting (at Boulderado Hotel)

18

JD291 (windy)
Site Visit (JB, MP, JH, NT, AW, Mike): Details

swapped 4 flasks, moved lowest inlet to top of green metal bar in the ground

very windy, ~zero degrees C, ~4 inches new snow.

Ameriflux Meeting (at Boulderado Hotel)

19

JD292
Ameriflux Meeting (at Boulderado Hotel)

20

JD293
Site Visit (DB): Details

21

JD294

22

JD295
Site Visit (JH, MP): Details

TIMES retreat (SpB off)

23

JD296
TIMES retreat (SpB off)

24

JD297 (warm, sunny)
Site Visit (JB, MP, JH): Details

downloaded from soil and Tcs

N2 @ 400 psi, span @ 700 psi

discovered the CR10/MRS logger box was open and full or ice/water!

25

JD298

26

JD299
Site Visit (JB, SpB, SO): Details , Photos (4)

lost about 3-4 hours of MET logger data on data system (this happend when we added a 2nd fiber optic converter to the quacker box for EOL/TRAM)

at ~9:30 MST, the re-furbished NCAR TC cr23x logger re-installed on tower

replaced the MRS cr10 logger with spare CR10 (from Jia).

from 13:30-14:20 MST, at soil logger---stopped saving the "tdr" parameters, started saving all 25 am25t mux TC data, started running "soil_02.csi"

27

JD300

28

JD301

29

JD302
Haze from CA fires present.

30

JD303
Site Visit (JB, SpB, JH, NT): Details , Photos (13)

rebooted isffa

downloaded from all loggers BUT cnr

replaced cable for the Rnet sensor at 2m on the mini tower; see details for more info

tower N2 @ 350 psi

at soil logger am25t mux---added 3 more TCs to the logger (now collecting data from all 25 TCs). setup the TCs near the snowtemp probes.

31

JD304
Halloween

1

JD305
Site Visit (JB, NT): Details

worked on last snow fence plot

swapped 4 flasks

2

JD306

3

JD307